OpenStudy (anonymous):

An inequality is shown below: −np − 5 ≤ 4(c − 2) Which of the following solves for n? n ≥ − the quantity 4 times c minus 3 all over p n ≥ − the quantity 4 times c minus 13 all over p n ≤ − the quantity 4 times c minus 3 all over p n ≤ − the quantity 4 times c minus 13 all over p

OpenStudy (anonymous):

@imqwerty

OpenStudy (calculusxy):

\[\large \rm \color {blue}{-np - 5 \le 4(c - 2)}\] \[\large \rm -np - 5\le 4c - 8\] \[\large \rm -np \le 4c -3 \] \[\large \rm -n \le \frac{ 4c - 3 }{ p }\] `switch the sign when it's negative like -n and change the sign` \[\large \rm n \ge \frac{ 4c - 3 }{ p }\]

OpenStudy (calculusxy):

I am not sure if this is exactly correct...

OpenStudy (anonymous):

@imqwerty what do u think?

imqwerty (imqwerty):

\[-np-5 \le 4(c-2) \]\[-np-5 \le 4c-8\]\[-np \le 4c-3\]dividing both sides by -p we will reverse the inequality cause we are dividing by a negative number sp\[n \ge \frac{ 4c-3 }{ -p }\]\[n \ge \frac{ -(4c-3) }{ p }\] oh no i didn't read calxy's solution :'( i typed it all its correct tho :')

OpenStudy (calculusxy):

why would it be negative "p"?

OpenStudy (calculusxy):

on the bottom?

imqwerty (imqwerty):

i divided both sides by -p :)

OpenStudy (calculusxy):

isn't n negative?

imqwerty (imqwerty):

no its not given anywhere

OpenStudy (calculusxy):

okay thanks :)

OpenStudy (anonymous):

what answer choice is that?

imqwerty (imqwerty):

no prblm B) it wuld be A

OpenStudy (anonymous):

Which graph represents the solutions to the inequality |2x − 4| less than or equal to 14? number line with an open circle on negative 5, shading to the right and an open circle on 9, shading to the left number line with a closed circle on negative 5, shading to the right and a closed circle on 9, shading to the left number line with an open circle on negative 5, shading to the left and an open circle on 9, shading to the right number line with a closed circle on negative 5, shading to the left and a closed circle on 9, shading to the right

OpenStudy (anonymous):

i think its D

OpenStudy (anonymous):

It is D

imqwerty (imqwerty):

\[|2x-4| \le 14\]\[-14 \le 2x-4 \le 14\]add 4 all over\[-10 \le 2x \le 18\]divide by 2\[-5 \le x \le 9\] so x can be between -5 and 9 so yes its D

OpenStudy (anonymous):

ok :D ty

imqwerty (imqwerty):

np :D